# Gravitational wave probes of particle dark matter: a review

### v4

@@ V.1 Annihilation signal
Ultralight bosons are their own anti-particles, which means that they continuously annihilate (even as the cloud is growing), producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1$, which predominately applies to the equations shown in this review). In the scalar boson case Abbott et al. (2022d):
@@
Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state. As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi et al. (2019):
@@
where we have explicitly noted a factor of 2 that is missing from the original reference Isi et al. (2019), as pointed out in Siemonsen et al. (2023b).
As the number of bosons starts to decrease, there will also be a classical depletion of the cloud over time. This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi et al. (2019) [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen et al. (2023b); May et al. (2025), which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
@@
In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi et al. (2019):
@@ V.2 Transition signal
In the scalar boson case, just as electrons jump between discrete energy levels in atoms, bosons can also transition between discrete energy states, provided these states are populated. The energy states of the boson cloud are quantized, with each state corresponding to specific quantum numbers. Significant transition signals are expected when the cloud reaches a population that allows transitions between particular energy levels, such as from the 5g to the 6g states, where “g” refers to the orbital angular momentum quantum number of the state. These transitions between specific energy levels are expected to generate detectable gravitational wave signals that could last for durations much longer than the typical observation times of ground-based gravitational wave interferometers. The frequency of these signals is given by Arvanitaki et al. (2015a):
@@
The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki et al. (2015a):
@@
While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki et al. (2015a), the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti et al. (2025).
In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen and East (2020). These higher overtones could saturate more quickly than the fundamental modes, allowing transitions to lower states and increasing the overall energy of the cloud, potentially leading to detectable gravitational wave emission.
@@ V.3 Impact of boson self-interactions on superradiance
    Saturation of Superradiance: Self-interactions modify the growth timescale, potentially leading to a faster saturation of the superradiance process. This could result in stronger GW emission once the cloud reaches its maximum energy state. The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino and Kodama (2012, 2015a, 2015b).
@@
    Merger Dynamics: The presence of self-interactions can affect how quickly the boson cloud dissipates or decays, influencing the coalescence of the black hole system. This decay modifies the timescale of the inspiral phase and the subsequent merger, altering the frequency evolution of the emitted GWs. A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon et al. (2024); Aurrekoetxea et al. (2024); Takahashi et al. (2024).
@@
    Enhanced Annihilation of Boson Clouds in Isolated Black Holes: In a non-interacting system, the annihilation rate of bosons is limited by the density and distribution of particles in the cloud. Self-interactions, however, can increase the density of the cloud by accelerating the process of boson accumulation and scattering. As a result, the rate of boson annihilation can be significantly enhanced. This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar et al. (2021).
@@
    Enhanced Energy-Level Transitions: The energy levels of bosons in the cloud are quantized, similar to the energy levels of electrons in an atom. Self-interactions can facilitate the transition of bosons between higher energy states more rapidly, increasing the amplitude of the emitted gravitational waves. This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar et al. (2021).
These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti et al. (2025). Recent searches for these signals have not accounted for these interactions, and their inclusion may shift current detection prospects by enhancing the signal strength or changing the emission characteristics. In this section, we will explore how self-interactions influence the dynamics of boson clouds and their implications for gravitational wave astronomy.”
Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated. The superradiant instability of spin-2 fields occurs much faster than the others East and Siemonsen (2023), and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen. Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra et al. (2024).
@@ V.4 Bosenova signal
If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda and Nakayama (2020); Collaviti et al. (2025); Omiya et al. (2024); Jones et al. (2025). The cloud may undergo a significant reconfiguration, leading to a dramatic release of energy, a process referred to as a “bosenova.” This occurs when the potential energy of the boson cloud becomes comparable to or exceeds the self-interaction energy, leading to the rapid rearrangement of the bosonic matter.
The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field Donley et al. (2001). However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar et al. (2021).
In particular, in the axion case, originally proposed in Arvanitaki and Dubovsky (2011), bosenovas would likely not occur. While early papers assumed that the boson cloud reached large enough amplitudes to produce GW emission Yoshino and Kodama (2012, 2015a, 2015b), subsequent studies in the non-relativistic Baryakhtar et al. (2021) and relativistic Omiya et al. (2023) regimes confirmed that dissipative nonlinear effects would prevent large-enough cloud amplitudes from occurring to cause a bosenova. In the vector case, bosenovas could actually occur if strings are formed the Higgs-Abelian mode East (2022); East and Huang (2022); Brzeminski et al. (2025).
Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time. In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki et al. (2010); Yoshino and Kodama (2012, 2015a, 2015b). This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered. Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, e.g. in Arvanitaki and Dubovsky (2011), permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie and Huang (2024) .
@@ V.5 Clouds in binary black hole systems
If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann et al. (2019a). In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann et al. (2019a, b, 2020) or through a dynamical friction drag force induced by the boson clouds. These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers Guo et al. (2024a). In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang and Yang (2019); Brito and Shah (2023), could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum. Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann et al. (2019a).
More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi et al. (2022); Baumann et al. (2022a, b). This ionization occurs when the energy from the binary black hole system becomes large enough to unbind the cloud from its host BH. As a result, the inspiral of the binary becomes significantly influenced by the dynamics of the boson cloud, rather than merely being perturbed by it.
@@ V.5.1 Spin-induced quadrupole moment
Ref. Krishnendu et al. (2017) and Ref. Rahman and Bhattacharyya (2023) showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen (1974), while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers and Poisson (1999); Pappas and Apostolatos (2012). In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
@@ V.5.2 Tidal disruptions
Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng et al. (2020). To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects. For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou (2020). For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann et al. (2019b). In Payne et al. (2022), the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers. Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes. Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds would significantly impact the formation and evolution histories of BHs.
@@ V.5.3 Binary-induced transitions
    Hyperfine Resonance: The orbital frequency of the binary system matches the energy spacing between two adjacent energy levels in the boson cloud. This condition is analogous to a resonance in atomic physics, where the energy difference between two states is exactly matched by an external driving frequency. In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann et al. (2019a).
@@
    Bohr Resonance: In addition to hyperfine resonance, the growing mode (the mode that is being excited by the inspiral) mixes with decaying modes of the cloud. This interaction allows energy to be transferred between different states, influencing the GW signal Baumann et al. (2019a).
@@
    Landau-Zener Transitions: When the orbital frequency of the inspiral matches the energy spacing between adjacent energy levels, the system undergoes a non-adiabatic shift in energy states. This transition process, characterized by the Landau-Zener mechanism, significantly alters the frequency and amplitude evolution of the GW signal. These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann et al. (2020).
These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann et al. (2019a).
@@ V.5.4 Ionization
When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi et al. (2022); Baumann et al. (2022a, b). The energy from the binary drives this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes dominated by the boson cloud, instead of simply perturbed by it. Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli et al. (2023). It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary. That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli et al. (2024).
Taken together, the various phenomena discussed above—from level transitions and resonances to the dramatic ionization of the cloud—highlight the rich interplay between boson clouds and binary black holes. Each mechanism offers a distinct pathway by which the presence of a boson cloud can alter binary evolution and leave characteristic imprints on the emitted GWs. These theoretical insights provide the foundation for future searches with space-based GW data Bertone et al. (2020); Bertone (2024). In the following section, however, we turn to current ground-based GW interferometer data, which so far constrain only annihilation signals from boson clouds, as discussed in Section V.1.

### v5

@@ V.1 Annihilation signal
Ultralight bosons are their own anti-particles, which means that they continuously annihilate (even as the cloud is growing), producing GW radiation at a frequency fixed by the boson mass and slightly redshifted by the BH mass (in the non-relativistic limit $\alpha\ll 1$, which predominately applies to the equations shown in this review). In the scalar boson case LIGOScientific:2021jlr:
@@
Qualitatively, the GWs emitted are monochromatic because the mass of the boson is fixed, and all the bosons are in the same energy state. As alluded to before, the timescales for GW emission are much longer than the time it takes to build the cloud Isi:2018pzk:
@@
where we have explicitly noted a factor of 2 that is missing from the original reference Isi:2018pzk, as pointed out in Siemonsen:2022yyf.
As the number of bosons starts to decrease, there will also be a classical depletion of the cloud over time. This reduction in bosons causes the cloud to contract, resulting in the following positive drifts in frequency over time Isi:2018pzk [^11] We note that relativistic computations of the frequency drift have also been performed Siemonsen:2022yyf; May:2024npn, which are useful for real searches of GWs arising from boson clouds but not less important for our discussions here.:
@@
In the non-relativistic limit $\alpha\ll 1$, the signals will have the corresponding amplitudes of Isi:2018pzk:
@@ V.2 Transition signal
In the scalar boson case, just as electrons jump between discrete energy levels in atoms, bosons can also transition between discrete energy states, provided these states are populated. The energy states of the boson cloud are quantized, with each state corresponding to specific quantum numbers. Significant transition signals are expected when the cloud reaches a population that allows transitions between particular energy levels, such as from the 5g to the 6g states, where “g” refers to the orbital angular momentum quantum number of the state. These transitions between specific energy levels are expected to generate detectable gravitational wave signals that could last for durations much longer than the typical observation times of ground-based gravitational wave interferometers. The frequency of these signals is given by Arvanitaki:2014wva:
@@
The frequency of the signal would drift over time according to the duration of the signal and the axion decay constant $f_{a}$ Arvanitaki:2014wva:
@@
While the expected event rate of these transitions in ground-based detectors is unlikely to exceed one event in around 3.5 years of observation with advanced LIGO Arvanitaki:2014wva, the presence of self-interactions could potentially enhance the transition rates, making these signals detectable Collaviti:2024mvh.
In the vector boson case, transitions can also occur, and interestingly, overtone modes may grow faster than the fundamental modes, particularly for small angular numbers, fine-structure constants, and black hole spin values Siemonsen:2019ebd. These higher overtones could saturate more quickly than the fundamental modes, allowing transitions to lower states and increasing the overall energy of the cloud, potentially leading to detectable gravitational wave emission.
@@ V.3 Impact of boson self-interactions on superradiance
    Saturation of Superradiance: Self-interactions modify the growth timescale, potentially leading to a faster saturation of the superradiance process. This could result in stronger GW emission once the cloud reaches its maximum energy state. The timescale for superradiance saturation is sensitive to the coupling strength between bosons, and self-interactions may allow the cloud to reach a critical size faster, enhancing the CW signal emitted during this phase Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa.
@@
    Merger Dynamics: The presence of self-interactions can affect how quickly the boson cloud dissipates or decays, influencing the coalescence of the black hole system. This decay modifies the timescale of the inspiral phase and the subsequent merger, altering the frequency evolution of the emitted GWs. A faster decay due to self-interactions could lead to a sharper rise in the frequency of the GW signal Boudon:2023vzl; Aurrekoetxea:2024cqd; Takahashi:2024fyq.
@@
    Enhanced Annihilation of Boson Clouds in Isolated Black Holes: In a non-interacting system, the annihilation rate of bosons is limited by the density and distribution of particles in the cloud. Self-interactions, however, can increase the density of the cloud by accelerating the process of boson accumulation and scattering. As a result, the rate of boson annihilation can be significantly enhanced. This increased annihilation rate leads to stronger and potentially more detectable GW signals from the cloud Baryakhtar:2020gao.
@@
    Enhanced Energy-Level Transitions: The energy levels of bosons in the cloud are quantized, similar to the energy levels of electrons in an atom. Self-interactions can facilitate the transition of bosons between higher energy states more rapidly, increasing the amplitude of the emitted gravitational waves. This could make energy-level transitions a more significant source of detectable GW signals, especially if the cloud is in a high-energy state Baryakhtar:2020gao.
These effects become particularly relevant when considering the potential detection of boson clouds in future GW observatories, such as Cosmic Explorer and ET Collaviti:2024mvh. Recent searches for these signals have not accounted for these interactions, and their inclusion may shift current detection prospects by enhancing the signal strength or changing the emission characteristics. In this section, we will explore how self-interactions influence the dynamics of boson clouds and their implications for gravitational wave astronomy.”
Note that while spin-0 and spin-1 boson clouds are relatively well understood theoretically, spin-2 clouds are much more complicated. The superradiant instability of spin-2 fields occurs much faster than the others East:2023nsk, and since the backreaction of the instability has not been computed due to difficulties arising from a nonlinear coupling between the spin-2 particle and gravity, it is not yet known whether the observational signatures are similar to the vector or scalar cases, and whether there is a dependency on the specific nonlinear theory chosen. Such ultralight bosons could also be used to probe the quantum nature of BHs Mitra:2023sny.
@@ V.4 Bosenova signal
If the self-interactions between bosons become stronger than the gravitational binding energy of the cloud, the cloud begins to destabilize and cannot be described analogously to the hydrogen atom Fukuda:2019ewf; Collaviti:2024mvh; Omiya:2024xlz; Jones:2024fpg. The cloud may undergo a significant reconfiguration, leading to a dramatic release of energy, a process referred to as a “bosenova.” This occurs when the potential energy of the boson cloud becomes comparable to or exceeds the self-interaction energy, leading to the rapid rearrangement of the bosonic matter.
The phenomenon of a bosenova has been observed in condensed matter systems, in which a Bose-Einstein condensate undergoes a similar process due to a change in interactions induced by an external magnetic field donley2001dynamics. However, the occurrence of a bosenova in the context of ultralight bosons around black holes remains highly uncertain, with studies suggesting that self-interactions may limit bosenovas to a restricted region of the allowed parameter space, particularly for certain combinations of black hole and boson masses Baryakhtar:2020gao.
In particular, in the axion case, originally proposed in Arvanitaki:2010sy, bosenovas would likely not occur. While early papers assumed that the boson cloud reached large enough amplitudes to produce GW emission Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa, subsequent studies in the non-relativistic Baryakhtar:2020gao and relativistic Omiya:2022gwu regimes confirmed that dissipative nonlinear effects would prevent large-enough cloud amplitudes from occurring to cause a bosenova. In the vector case, bosenovas could actually occur if strings are formed the Higgs-Abelian mode East:2022ppo; East:2022rsi; Brzeminski:2024drp.
Though GWs from boson clouds described above may seem independent, they could in principle be happening at the same time. In particular, if ultralight scalar or vector bosons have some self-interactions, and If these clouds are composed of string axions (as opposed to the QCD axion), a non-linear self-interaction may cause a bosenova if, during superradiance, the value of the scalar field approaches the axion decay constant $f_{a}$, which could be below the grand unification theory (GUT) scale Arvanitaki:2009fg; Yoshino:2012kn; Yoshino:2014wwa; Yoshino:2015nsa. This bosenova would halt the superradiance process, as about 5% of the boson cloud mass would fall back into the BH, but superradiance would immediately start again until the next bosenova is triggered. Such periodic bosenovas could sustain superradiance for much longer durations than those in the conventional scenario, e.g. in Arvanitaki:2010sy, permitting annihilating clouds even around extremely old BHs, and even affect GWs from binary black hole systems Xie:2022uvp .
@@ V.5 Clouds in binary black hole systems
If boson clouds form around binary black hole systems, they should imprint some signature in the GWs arising from their inspiral Baumann:2018vus. In particular, the GW signal will be modified due to tidal disruptions induced by the companion, the multiple moments of the cloud Baumann:2018vus; Baumann:2019eav; Baumann:2019ztm or through a dynamical friction drag force induced by the boson clouds. These kinds of signals could be visible in future space-based detectors, since the deviations primarily occur in the early inspiral stages of the system, i.e. at frequencies much lower than currently accessible by ground-based GW interferometers Guo:2023lbv. In both cases, observations of the inspiral of binary systems, and in particular EMRI systems Zhang:2018kib; Brito:2023pyl, could indicate the presence of boson clouds around one or both of the objects if there is some dephasing with respect to the waveform in vacuum. Additionally, once the binary reaches a certain stage of its evolution, the impact of both the spin-induced quadrupole moment and tidal forces will lessen, since the clouds will become disrupted by their companions. If we can observe variations in both the spin-induced quadrupole moment and the tidal deformability over the course of the lifetime of the binary, it would indicate the presence of a (disrupted) boson cloud Baumann:2018vus.
More recently, it has been shown that as the orbital separation in a binary black hole system decreases and becomes comparable to the size of the boson cloud, the boson cloud could be “ionized” Takahashi:2021yhy; Baumann:2021fkf; Baumann:2022pkl. This ionization occurs when the energy from the binary black hole system becomes large enough to unbind the cloud from its host BH. As a result, the inspiral of the binary becomes significantly influenced by the dynamics of the boson cloud, rather than merely being perturbed by it.
@@ V.5.1 Spin-induced quadrupole moment
Ref. Krishnendu:2017shb and Ref. Rahman:2021eay showed that $\kappa$ can be used as a distinguishing parameter for different kinds of objects, and also a test of the no-hair theorem: as an example, $\kappa=1$ for BHs Hansen:1974zz, while $\kappa\sim[1.4,8]$ for neutron stars depending on the equation-of-state Laarakkers:1997hb; Pappas:2012ns. In the case of boson cloud systems, $\kappa$ will take on vastly different values, of $\mathcal{O}(1000)$, depending on the boson cloud and BH masses, and the extent to which the Kerr metric is altered in the presence of a cloud with enough mass compared to the BH.
@@ V.5.2 Tidal disruptions
Because the boson cloud is much less compact than the BH, we expect that tidal disruptions from the companion could occur when the two objects in the binary black hole are close enough to each other Ng:2020jqd. To characterize the effects of tides in a binary system, the tidal deformability parameter $\Lambda$ is used, which depends on the mass and radius of the compact objects. For BHs in vacuum, this quantity is 0, while for neutron stars, it can take on values of $\mathcal{O}(100-1000)$ Chatziioannou:2020pqz. For boson clouds, $\Lambda\sim 10^{7}$, which is a remarkable departure from that expected from conventional astrophysical binaries Baumann:2019eav. In Payne:2021ahy, the authors considered the impact that superradiance would have on hierarchical BH mergers, and showed that the reduction of spin in BHs implies smaller recoil velocities of binary black hole mergers. Smaller recoil velocities imply that remnants could remain inside clusters, leading to higher chances of hierarchical formation and merger of binary black holes. Such disruptive effects of bosons on not just individual BHs but also on the BH population, imply that boson clouds would significantly impact the formation and evolution histories of BHs.
@@ V.5.3 Binary-induced transitions
    Hyperfine Resonance: The orbital frequency of the binary system matches the energy spacing between two adjacent energy levels in the boson cloud. This condition is analogous to a resonance in atomic physics, where the energy difference between two states is exactly matched by an external driving frequency. In this case, the system stays in equilibrium, and energy is transferred between energy levels that are in resonance Baumann:2018vus.
@@
    Bohr Resonance: In addition to hyperfine resonance, the growing mode (the mode that is being excited by the inspiral) mixes with decaying modes of the cloud. This interaction allows energy to be transferred between different states, influencing the GW signal Baumann:2018vus.
@@
    Landau-Zener Transitions: When the orbital frequency of the inspiral matches the energy spacing between adjacent energy levels, the system undergoes a non-adiabatic shift in energy states. This transition process, characterized by the Landau-Zener mechanism, significantly alters the frequency and amplitude evolution of the GW signal. These non-adiabatic transitions can provide observable changes in the signal, offering additional insights into the properties of the boson cloud Baumann:2019ztm.
These effects, all of which could be present simultaneously, would provide “smoking-gun” signatures of the presence of boson clouds around rotating black holes in a binary system Baumann:2018vus.
@@ V.5.4 Ionization
When the binary system separation becomes comparable to the size of the cloud, an “ionization” of the boson cloud occurs as the cloud becomes unbounded from its host BH Takahashi:2021yhy; Baumann:2021fkf; Baumann:2022pkl. The energy from the binary drives this effect, and is significantly larger than that emitted via GWs; thus, the inspiral of the binary becomes dominated by the boson cloud, instead of simply perturbed by it. Ionization will also tend to circularize the binary if it formed through dynamical capture, while leaving any orbital inclinations unaffected Tomaselli:2023ysb. It has been recently found, however, that as the resonances in the boson cloud evolve during the orbit, the boson cloud is unperturbed and ordinarily visible in LISA if the cloud and binary are counter rotating; otherwise, the cloud is destroyed because of the resonances, but a distinct mark is left on the binary. That is, the binary is forced to co-rotate with the boson cloud, and its eccentricity is driven to a particular value, which allows the possibility of doing statistics with detected binary black hole systems in LISA to probe the boson cloud hypothesis Tomaselli:2024dbw.
Taken together, the various phenomena discussed above—from level transitions and resonances to the dramatic ionization of the cloud—highlight the rich interplay between boson clouds and binary black holes. Each mechanism offers a distinct pathway by which the presence of a boson cloud can alter binary evolution and leave characteristic imprints on the emitted GWs. These theoretical insights provide the foundation for future searches with space-based GW data Bertone:2019irm; Bertone:2024rxe. In the following section, however, we turn to current ground-based GW interferometer data, which so far constrain only annihilation signals from boson clouds, as discussed in Section V.1.